The product of my number and twice my number is 128. What is half my number? Explain how you solved the problem. 4th grade math.
Answer by ReadingBoosters(3246)

x is the number
"product of" is multiplication
(x)(2x)=128
Distribute the x
2x^2 = 128
Divide both sides by 2
x^2 = 64
squareroot(x^2) = squareroot(64)
x = 8
Half of x is x/2
8/2 = 4 as half the number
.......................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.